### v1

# Daniele Bartoli1, Mohit Pal2, Pantelimon Stănică3, Tommaso Toccotelli1

Abstract
In this note we show (for a large enough dimension of the underlying field) a conjecture of [C. Beierle, C. Carlet, G. Leander, L. Perrin, A further study of quadratic APN permutations in dimension nine, Finite Fields Appl. 81 (2022), 102049] on a trivariate permutation. This function is a global representation of two new sporadic quadratic APN permutations in dimension $9$ found by [C. Beierle, G. Leander, New instances of quadratic APN functions, IEEE Trans. Inf. Theory 68(1) (2022), 670–678].
Keywords: Finite fields, Permutation polynomials, Differential uniformity, varieties, irreducible components
Mathematics Subject Classification 2020: 12E20, 11T06
1 Introduction and tools from algebraic geometry
Let $q=2^{m},$ $m\in\mathbb{N},$ and denote by $\mathbb{F}_{q}$ the finite field with $q$ elements. For any positive integer $n$, we denote by $\mathbb{F}_{q}[X_{1},\ldots,X_{n}]$, the ring of polynomials in $n$ indeterminates over finite field $\mathbb{F}_{q}$. An element $f\in\mathbb{F}_{q}[X_{1},\ldots,X_{n}]$ is called a permutation polynomial in $n$ variables if the equation $f(X_{1},\ldots X_{n})=a$ has $q^{n-1}$ solutions in $\mathbb{F}_{q}^{n}$ for each $a\in\mathbb{F}_{q}$. Let $F:\mathbb{F}_{q}^{n}\rightarrow\mathbb{F}_{q}^{n}$ be a map given by
$F(X_{1},\ldots,X_{n})=(f_{1}(X_{1},\ldots,X_{n}),\ldots,f_{n}(X_{1},\ldots,X_{n})),$
where $f_{i}\in\mathbb{F}_{q}[X_{1},\ldots,X_{n}]$, then $F$ is called a vectorial permutation if it induces a permutation on $\mathbb{F}_{q}^{n}$.
Vectorial Boolean functions are fundamental building blocks in symmetric cryptography, since many block ciphers employ these as components in their S-boxes. The security of a block cipher depends upon the immunity of its substitution boxes against various kinds of cryptographic attacks. For instance, a low differential uniformity [13] to resist the differential attacks [5]. For a prime $p$ and positive integer $n>0$, the differential uniformity of an $(n,n)$-function $F:\mathbb{F}_{p^{n}}\to\mathbb{F}_{p^{n}}$ is defined as the maximum number of solutions $x$ of the differential equation $F(x+a)-F(x)=b$, where $a\neq 0,b\in\mathbb{F}_{p^{n}}$. When then differential uniformity is $1$, we say that the function is perfect nonlinear (or planar). It may be noted that planar functions exist only in the case of odd characteristic. The lowest possible differential uniformity of functions over finite fields of even characteristic is $2$ and such functions are called almost perfect nonlinear (APN). APN functions in characteristic $2$ have been studied since the 90’s and a lot of research has been done in this direction in recent years (see [7] and references therein).
For the necessary algebraic geometry background, we refer to [9, 10]. We use the notations $\mathbb{P}^{r}(\mathbb{F}_{q})$ and $\mathbb{A}^{r}(\mathbb{F}_{q})$ (or $\mathbb{F}_{q}^{r}$) to denote the projective and the affine space of dimension $r\in\mathbb{N}$ over the finite field $\mathbb{F}_{q}$, respectively. A variety, and more specifically a curve or a surface (i.e. a variety of dimension 1 or 2, respectively), is described by a certain set of equations with coefficients in $\mathbb{F}_{q}.$ We say that a variety $\mathcal{V}$ is absolutely irreducible if there are no varieties $\mathcal{V}^{\prime}$ and $\mathcal{V}^{\prime\prime}$ defined over the algebraic closure of $\mathbb{F}_{q}$ (denoted by $\overline{\mathbb{F}_{q}}$) and different from $\mathcal{V}$ such that $\mathcal{V}=\mathcal{V}^{\prime}\cup\mathcal{V}^{\prime\prime}$. If a variety $\mathcal{V}\subseteq\mathbb{A}^{r}(\mathbb{F}_{q})$ is defined by $F_{i}(X_{1},\dots,X_{r})=0,$ for $i=1,\dots s,$ an $\mathbb{F}_{q}$-rational point of $\mathcal{V}$ is a point $(x_{1},\dots,x_{r})\in\mathbb{A}^{r}(\mathbb{F}_{q})$ such that $F_{i}(x_{1},\dots,x_{r})=0,$ for $i=1,\dots,s.$ The set of the $\mathbb{F}_{q}$-rational points of $\mathcal{V}$ is usually denoted by $\mathcal{V}(\mathbb{F}_{q}).$ If $s=1,$ $\mathcal{V}$ is called a hypersurface and it is absolutely irreducible if the corresponding polynomial $F(X_{1},\dots,X_{r})$ is absolutely irreducible, i.e. it possesses no non-constant factors over $\overline{\mathbb{F}_{q}}.$ Moreover, we say that $\mathcal{V}$ is a variety of degree $d$ (and write $\deg(\mathcal{V})=d$) if $d=\#(\mathcal{V}\cap H),$ where $H\subseteq\mathbb{A}^{r}(\overline{\mathbb{F}_{q}})$ is a general projective subspace of dimension $r-s.$ To determine the degree of a variety is generally not straightforward; however an upper bound to $\deg(\mathcal{V)}$ is given by $\prod_{i=1}^{s}\deg(F_{i}).$ We also recall that the Frobenius map $\Phi_{q}:x\mapsto x^{q}$ is an automorphism of $\mathbb{F}_{q^{k}}$ and generates the group $Gal(\mathbb{F}_{q^{k}}/\mathbb{F}_{q})$ of automorphisms of $\mathbb{F}_{q^{k}}$ that fixes $\mathbb{F}_{q}$ pointwise. The Frobenius automorphism induces also a collineation of $\mathbb{A}^{r}(\overline{\mathbb{F}_{q}})$ and an automorphism of $\overline{\mathbb{F}_{q}}[X_{1},\dots,X_{r}].$
A crucial point in our investigation of permutation trinomials over $\mathbb{F}_{q^{3}}$ is to prove the existence of suitable $\mathbb{F}_{q}$-rational points in algebraic surfaces $\mathcal{V}$ attached to each permutation trinomial. This is reached by proving the existence of absolutely irreducible $\mathbb{F}_{q}$-rational components in $\mathcal{V}$ and lower bounding the number of their $\mathbb{F}_{q}$-rational points. To this end, generalizations of Hasse-Weil type bounds for algebraic varieties are needed. To ensure the existence of a suitable $\mathbb{F}_{q}$-rational point of $\mathcal{V},$ we need the following result.
Theorem 1.
[8, Theorem 7.1] Let $\mathcal{V}\subseteq\mathbb{A}^{n}(\mathbb{F}_{q})$ be an absolutely irreducible variety defined over $\mathbb{F}_{q}$ of dimension $r>0$ and degree $\delta.$ If $q>2(r+1)\delta^{2},$ then the following estimate holds:
|#$\big{|}\#(\mathcal{V}(\mathbb{A}^{n}(\mathbb{F}_{q})))-q^{r}\big{|}\leq(\delta-1)(\delta-2)q^{r-1/2}+5\delta^{13/3}q^{r-1}.$ (1)
In our approach we will make use of the following result.
Lemma 2.
[1, Lemma 2.1] Let $\mathcal{H}$ be a projective hypersurface and $\mathcal{X}$ a projective variety of dimension $n-1$ in $\mathbb{P}^{n}(\mathbb{F}_{q})$. If $\mathcal{X}\cap\mathcal{H}$ has a non-repeated absolutely irreducible component defined over $\mathbb{F}_{q}$ then $\mathcal{X}$ has a non-repeated absolutely irreducible component defined over $\mathbb{F}_{q}$.
2 Main result
Beierle and Leander [4] found two new APN permutations in dimension $9$, namely
$x\mapsto x^{3}+ux^{10}+u^{2}x^{17}+u^{4}x^{80}+u^{5}x^{192},$
$x\mapsto x^{3}+u^{2}x^{10}+ux^{24}+u^{4}x^{80}+u^{6}x^{136}.$
In this note, we shall consider the trivariate function $C_{u}:\mathbb{F}_{q}^{3}\rightarrow\mathbb{F}_{q}^{3}$ given by
$(X,Y,Z)\mapsto(X^{3}+uY^{2}Z,Y^{3}+uXZ^{2},Z^{3}+uX^{2}Y).$
This trivariate function is the representation of the two APN permutations in dimension $9$ found by [4] mentioned above, and has been considered in [3] where the authors have shown that it is a vectorial permutation for $m=3$ and $u\in\mathbb{F}_{2^{m}}\backslash\{0,1\}$. For $m>3$, the authors have proposed the following conjecture.
Conjecture 3 ([3]).
Let $m>3$ and let $u\in\mathbb{F}_{2^{m}}^{*}$. Then the function $C_{u}:\mathbb{F}_{2^{m}}^{3}\to\mathbb{F}_{2^{m}}^{3}$ given by
$(X,Y,Z)\mapsto(X^{3}+uY^{2}Z,Y^{3}+uXZ^{2},Z^{3}+uX^{2}Y)$
is not a permutation, nor APN.
In 2022, Bartoli and Timpanella [2] proved that $C_{u}$ is not APN for $m\geq 20$. Together with the experimental results in small dimensions, this proved the second part of the above conjecture, for $m\geq 3$. It is the intent of this note to prove the remaining part of the above conjecture. It is worth mentioning here that the APN permutations given in [4] are generalised into an infinite family of APN functions in trivariate form in [12].
It is easy to observe that when $m$ is even then $C_{u}$ is not a permutation as in this case $C_{u}(X,0,0)=(X^{3},0,0)$ and the function $X\mapsto X^{3}$ is $3$-to-$1$. In what follows we assume that $m$ is odd. Note that $C_{u}$ is a permutation if and only if the equation
$C_{u}(X+\alpha,Y+\beta,Z+\gamma)+C_{u}(X,Y,Z)=0$
has only the trivial solutions $\{(x,y,z,0,0,0)\ :\ x,y,z\in\mathbb{F}_{2^{m}}\}$.
Such a condition reads
$\begin{cases}\alpha X^{2}+\alpha^{2}X+u\gamma Y^{2}+u\beta^{2}Z&=\alpha^{3}+u\beta^{2}\gamma\\ \beta Y^{2}+\beta^{2}Y+u\gamma^{2}X+u\alpha Z^{2}&=\beta^{3}+u\gamma^{2}\alpha\\ \gamma Z^{2}+\gamma^{2}Z+u\beta X^{2}+u\alpha^{2}Y&=\gamma^{3}+u\alpha^{2}\beta.\end{cases}$ (2)
Before we take our algebraic geometry approach, we make some observations. First, we assume that only one among $\alpha,\beta$ and $\gamma$ is nonzero. Without loss of generality (here because of the symmetric property of $C_{u}$), we may assume that $\alpha\neq 0$ and $\beta=\gamma=0$. Then System (2) becomes
$\begin{cases}\alpha X^{2}+\alpha^{2}X+\alpha^{3}&=0\\ u\alpha Z^{2}&=0\\ u\alpha^{2}Y&=0.\end{cases}$
It is straightforward to see that the first equation of the above system has no solution $(\overline{x},\overline{y},\overline{z})\in\mathbb{F}_{2^{m}}^{3}$ as $m$ is odd.
Next, we assume that only one among $\alpha,\beta$ and $\gamma$ is zero. Again, we may assume that $\alpha,\beta\neq 0$ and $\gamma=0$. Then System (2) becomes
$\begin{cases}\alpha X^{2}+\alpha^{2}X+u\beta^{2}Z&=\alpha^{3}\\ \beta Y^{2}+\beta^{2}Y+u\alpha Z^{2}&=\beta^{3}\\ u\beta X^{2}+u\alpha^{2}Y&=u\alpha^{2}\beta,\end{cases}$ (3)
which can be further simplified by replacing $X\mapsto\alpha X$ and $Y\mapsto\beta Y$, i.e., System (3) is equivalent to the following system
$\begin{cases}\alpha^{3}(X^{2}+X+1)+u\beta^{2}Z&=0\\ \beta^{3}(Y^{2}+Y+1)+u\alpha Z^{2}&=0\\ X^{2}+Y+1&=0.\end{cases}$ (4)
Now, squaring the first equation and putting $X^{2}=Y+1$ into it, we have
$\alpha^{6}(Y^{2}+Y+1)+u^{2}\beta^{4}Z^{2}=0.$
Multiplying above equation by $\alpha$ and putting $\ u\alpha Z^{2}=\beta^{3}(Y^{2}+Y+1)$, we have
$(\alpha^{7}+u\beta^{7})(Y^{2}+Y+1)=0.$
Thus, System (4) is equivalent to the following system
$\begin{cases}(\alpha^{7}+u\beta^{7})(Y^{2}+Y+1)&=0\\ \beta^{3}(Y^{2}+Y+1)+u\alpha Z^{2}&=0\\ X^{2}+Y+1&=0.\end{cases}$
Notice that if $\alpha^{7}+u\beta^{7}=0$, i.e., $u$ is a $7^{th}$ power then the above system has nonzero solutions $(\overline{x},\overline{y},\overline{z})\in\mathbb{F}_{2^{m}}^{3}$ and consequently $C_{u}$ is not a permutation. When $u$ is not a $7^{th}$ power then the first equation of this system has no solution $\overline{y}\in\mathbb{F}_{2^{m}}$ for $m$ odd. Thus, in what follows, we assume that $u$ is not a $7^{th}$ power and $m$ is odd.
Let
$f(\alpha,\beta,X,Y,Z)$ $:=\frac{\alpha^{3}+\alpha^{2}X+\alpha X^{2}+\beta^{2}Zu}{\beta^{2}u+Y^{2}u},$
$g(\beta,X,Y,Z)$ $:=\beta^{12}u^{9}+\beta^{12}u^{6}+\beta^{12}u^{3}+\beta^{12}+\beta^{9}XZ^{2}u^{10}+\beta^{9}XZ^{2}u^{4}+\beta^{9}Y^{3}u^{9}+\beta^{9}Y^{3}u^{3}$
$+\beta^{8}XYZ^{2}u^{10}+\beta^{8}XYZ^{2}u^{4}+\beta^{8}Y^{4}u^{6}+\beta^{8}Y^{4}+\beta^{6}X^{2}Z^{4}u^{8}+\beta^{6}X^{2}Z^{4}u^{5}$
$+\beta^{6}Y^{6}u^{6}+\beta^{6}Y^{6}u^{3}+\beta^{5}X^{7}u^{8}+\beta^{5}X^{7}u^{5}+\beta^{5}XY^{4}Z^{2}u^{10}+\beta^{5}XY^{4}Z^{2}u^{7}$
$+\beta^{4}X^{7}Yu^{8}+\beta^{4}X^{7}Yu^{5}+\beta^{4}X^{2}Y^{2}Z^{4}u^{8}+\beta^{4}X^{2}Y^{2}Z^{4}u^{5}+\beta^{4}XY^{5}Z^{2}u^{10}$
$+\beta^{4}XY^{5}Z^{2}u^{7}+\beta^{4}Y^{8}u^{9}+\beta^{4}Y^{8}+\beta^{3}X^{7}Y^{2}u^{5}+\beta^{3}X^{4}Y^{4}Zu^{6}+\beta^{3}X^{3}Z^{6}u^{6}$
$+\beta^{3}X^{3}Z^{6}u^{3}+\beta^{3}X^{2}Y^{3}Z^{4}u^{5}+\beta^{3}XY^{6}Z^{2}u^{4}+\beta^{3}Y^{9}u^{3}+\beta^{3}Y^{2}Z^{7}u^{4}$
$+\beta^{2}X^{8}Z^{2}u^{6}+\beta^{2}X^{7}Y^{3}u^{8}+\beta^{2}X^{5}Y^{2}Z^{3}u^{7}+\beta^{2}X^{4}Y^{5}Zu^{6}+\beta^{2}X^{3}YZ^{6}u^{3}$
$+\beta^{2}XY^{7}Z^{2}u^{7}+\beta^{2}XY^{7}Z^{2}u^{4}+\beta^{2}XZ^{9}u^{5}+\beta^{2}Y^{10}u^{6}+\beta^{2}Y^{3}Z^{7}u^{4}$
$+\beta X^{7}Y^{4}u^{8}+\beta X^{4}Y^{6}Zu^{6}+\beta X^{3}Y^{2}Z^{6}u^{6}+\beta X^{3}Y^{2}Z^{6}u^{3}+\beta X^{2}Y^{5}Z^{4}u^{5}$
$+\beta XY^{8}Z^{2}u^{7}+\beta Y^{11}u^{9}+\beta Y^{4}Z^{7}u^{4},$
$h(\beta,X,Y,Z)$ $:=u^{5}(\beta+Y)^{2}(\beta^{3}X^{6}u^{3}+\beta^{3}X^{6}+\beta^{3}Y^{4}Z^{2}u^{5}+\beta^{3}Y^{4}Z^{2}u^{2}+\beta^{2}X^{6}Yu^{3}+\beta^{2}X^{6}Y$
$+\beta^{2}Y^{5}Z^{2}u^{5}+\beta^{2}Y^{5}Z^{2}u^{2}+\beta X^{6}Y^{2}u^{3}+\beta X^{6}Y^{2}+\beta Y^{6}Z^{2}u^{5}+\beta Y^{6}Z^{2}u^{2}$
$+X^{7}Z^{2}u+X^{4}Y^{2}Z^{3}u^{2}+X^{2}YZ^{6}u+XY^{4}Z^{4}u^{3}+Y^{7}Z^{2}u^{5}+Z^{9}),$
$a_{0}(\beta,Y,Z)$ $:=\beta^{24}u^{21}+\beta^{24}u^{18}+\beta^{24}u^{15}+\beta^{24}u^{12}+\beta^{24}u^{9}+\beta^{24}u^{6}+\beta^{24}u^{3}+\beta^{24}$
$+\beta^{21}Y^{3}u^{21}+\beta^{21}Y^{3}u^{15}+\beta^{21}Y^{3}u^{9}+\beta^{21}Y^{3}u^{3}+\beta^{20}Y^{4}u^{21}+\beta^{20}Y^{4}u^{15}$
$+\beta^{20}Y^{4}u^{9}+\beta^{20}Y^{4}u^{3}+\beta^{18}Y^{6}u^{18}+\beta^{18}Y^{6}u^{15}+\beta^{18}Y^{6}u^{6}+\beta^{18}Y^{6}u^{3}$
$+\beta^{17}Y^{7}u^{21}+\beta^{17}Y^{7}u^{15}+\beta^{17}Y^{7}u^{9}+\beta^{17}Y^{7}u^{3}+\beta^{16}Y^{8}u^{15}+\beta^{16}Y^{8}u^{12}$
$+\beta^{16}Y^{8}u^{3}+\beta^{16}Y^{8}+\beta^{15}Y^{9}u^{15}+\beta^{15}Y^{9}u^{3}+\beta^{15}Y^{2}Z^{7}u^{16}+\beta^{15}Y^{2}Z^{7}u^{4}$
$+\beta^{14}Y^{10}u^{15}+\beta^{14}Y^{10}u^{3}+\beta^{14}Y^{3}Z^{7}u^{16}+\beta^{14}Y^{3}Z^{7}u^{4}+\beta^{13}Y^{11}u^{15}$
$+\beta^{13}Y^{11}u^{3}+\beta^{13}Y^{4}Z^{7}u^{16}+\beta^{13}Y^{4}Z^{7}u^{4}+\beta^{12}Y^{12}u^{15}+\beta^{12}Y^{12}u^{12}$
$+\beta^{12}Y^{12}u^{9}+\beta^{12}Y^{12}u^{6}+\beta^{11}Y^{13}u^{15}+\beta^{11}Y^{13}u^{3}+\beta^{11}Y^{6}Z^{7}u^{16}$
$+\beta^{11}Y^{6}Z^{7}u^{4}+\beta^{10}Y^{14}u^{15}+\beta^{10}Y^{14}u^{3}+\beta^{10}Y^{7}Z^{7}u^{16}+\beta^{10}Y^{7}Z^{7}u^{4}$
$+\beta^{9}Y^{15}u^{15}+\beta^{9}Y^{15}u^{9}+\beta^{9}Y^{8}Z^{7}u^{22}+\beta^{9}Y^{8}Z^{7}u^{4}+\beta^{8}Y^{16}u^{21}+\beta^{8}Y^{16}u^{18}$
$+\beta^{8}Y^{16}u^{9}+\beta^{8}Y^{16}+\beta^{8}Y^{9}Z^{7}u^{22}+\beta^{8}Y^{9}Z^{7}u^{16}+\beta^{7}Y^{17}u^{15}+\beta^{7}Y^{17}u^{3}$
$+\beta^{7}Y^{10}Z^{7}u^{16}+\beta^{7}Y^{10}Z^{7}u^{4}+\beta^{6}Y^{18}u^{15}+\beta^{6}Y^{18}u^{6}+\beta^{6}Y^{11}Z^{7}u^{22}$
$+\beta^{6}Y^{11}Z^{7}u^{4}+\beta^{6}Y^{4}Z^{14}u^{11}+\beta^{6}Y^{4}Z^{14}u^{8}+\beta^{5}Y^{19}u^{21}+\beta^{5}Y^{19}u^{9}$
$+\beta^{5}Y^{12}Z^{7}u^{16}+\beta^{5}Y^{12}Z^{7}u^{4}+\beta^{4}Y^{20}u^{21}+\beta^{4}Y^{20}u^{12}+\beta^{4}Y^{13}Z^{7}u^{22}$
$+\beta^{4}Y^{13}Z^{7}u^{16}+\beta^{4}Y^{6}Z^{14}u^{11}+\beta^{4}Y^{6}Z^{14}u^{8}+\beta^{3}Y^{21}u^{15}+\beta^{3}Y^{14}Z^{7}u^{16}$
$+\beta^{3}Y^{7}Z^{14}u^{11}+\beta^{3}Z^{21}u^{12}+\beta^{2}Y^{22}u^{18}+\beta^{2}Y^{15}Z^{7}u^{22}+\beta^{2}Y^{8}Z^{14}u^{8}$
$+\beta^{2}YZ^{21}u^{12}+\beta Y^{23}u^{21}+\beta Y^{16}Z^{7}u^{22}+\beta Y^{9}Z^{14}u^{11}+\beta Y^{2}Z^{21}u^{12},$
$a_{1}(\beta,Y,Z)$ $:=\beta^{21}Z^{2}u^{19}+\beta^{21}Z^{2}u^{13}+\beta^{21}Z^{2}u^{7}+\beta^{21}Z^{2}u+\beta^{20}YZ^{2}u^{19}+\beta^{20}YZ^{2}u^{13}$
$+\beta^{20}YZ^{2}u^{7}+\beta^{20}YZ^{2}u+\beta^{18}Y^{3}Z^{2}u^{19}+\beta^{18}Y^{3}Z^{2}u^{13}+\beta^{18}Y^{3}Z^{2}u^{7}$
$+\beta^{18}Y^{3}Z^{2}u+\beta^{17}Y^{4}Z^{2}u^{19}+\beta^{17}Y^{4}Z^{2}u^{13}+\beta^{17}Y^{4}Z^{2}u^{7}+\beta^{17}Y^{4}Z^{2}u$
$+\beta^{15}Y^{6}Z^{2}u^{13}+\beta^{15}Y^{6}Z^{2}u+\beta^{14}Y^{7}Z^{2}u^{13}+\beta^{14}Y^{7}Z^{2}u+\beta^{13}Y^{8}Z^{2}u^{13}$
$+\beta^{13}Y^{8}Z^{2}u+\beta^{11}Y^{10}Z^{2}u^{13}+\beta^{11}Y^{10}Z^{2}u+\beta^{10}Y^{11}Z^{2}u^{13}+\beta^{10}Y^{11}Z^{2}u$
$+\beta^{9}Y^{12}Z^{2}u^{13}+\beta^{9}Y^{12}Z^{2}u^{7}+\beta^{8}Y^{13}Z^{2}u^{7}+\beta^{8}Y^{13}Z^{2}u+\beta^{7}Y^{14}Z^{2}u^{13}$
$+\beta^{7}Y^{14}Z^{2}u+\beta^{6}Y^{15}Z^{2}u^{13}+\beta^{6}Y^{15}Z^{2}u^{7}+\beta^{5}Y^{16}Z^{2}u^{19}+\beta^{5}Y^{16}Z^{2}u^{7}$
$+\beta^{4}Y^{17}Z^{2}u^{19}+\beta^{4}Y^{17}Z^{2}u^{13}+\beta^{3}Y^{18}Z^{2}u^{13}+\beta^{3}Y^{4}Z^{16}u^{9}+\beta^{2}Y^{19}Z^{2}u^{19}$
$+\beta^{2}Y^{5}Z^{16}u^{9}+\beta Y^{20}Z^{2}u^{19}+\beta Y^{6}Z^{16}u^{9},$
$a_{2}(\beta,Y,Z)$ $:=\beta^{18}Z^{4}u^{23}+\beta^{18}Z^{4}u^{20}+\beta^{18}Z^{4}u^{11}+\beta^{18}Z^{4}u^{8}+\beta^{16}Y^{2}Z^{4}u^{23}$
$+\beta^{16}Y^{2}Z^{4}u^{20}+\beta^{16}Y^{2}Z^{4}u^{11}+\beta^{16}Y^{2}Z^{4}u^{8}+\beta^{15}Y^{3}Z^{4}u^{23}+\beta^{15}Y^{3}Z^{4}u^{11}$
$+\beta^{14}Y^{4}Z^{4}u^{23}+\beta^{14}Y^{4}Z^{4}u^{11}+\beta^{13}Y^{5}Z^{4}u^{23}+\beta^{13}Y^{5}Z^{4}u^{11}+\beta^{11}Y^{7}Z^{4}u^{23}$
$+\beta^{11}Y^{7}Z^{4}u^{11}+\beta^{10}Y^{8}Z^{4}u^{23}+\beta^{10}Y^{8}Z^{4}u^{11}+\beta^{9}Y^{9}Z^{4}u^{23}+\beta^{9}Y^{9}Z^{4}u^{17}$
$+\beta^{9}Y^{2}Z^{11}u^{18}+\beta^{9}Y^{2}Z^{11}u^{6}+\beta^{8}Y^{10}Z^{4}u^{17}+\beta^{8}Y^{10}Z^{4}u^{11}+\beta^{8}Y^{3}Z^{11}u^{18}$
$+\beta^{8}Y^{3}Z^{11}u^{6}+\beta^{7}Y^{11}Z^{4}u^{23}+\beta^{7}Y^{11}Z^{4}u^{11}+\beta^{6}Y^{12}Z^{4}u^{23}+\beta^{6}Y^{12}Z^{4}u^{17}$
$+\beta^{6}Y^{12}Z^{4}u^{11}+\beta^{6}Y^{12}Z^{4}u^{8}+\beta^{6}Y^{5}Z^{11}u^{18}+\beta^{6}Y^{5}Z^{11}u^{6}+\beta^{5}Y^{13}Z^{4}u^{23}$
$+\beta^{5}Y^{13}Z^{4}u^{11}+\beta^{4}Y^{14}Z^{4}u^{17}+\beta^{4}Y^{14}Z^{4}u^{8}+\beta^{4}Y^{7}Z^{11}u^{18}+\beta^{4}Y^{7}Z^{11}u^{6}$
$+\beta^{3}Y^{15}Z^{4}u^{23}+\beta^{3}Y^{15}Z^{4}u^{11}+\beta^{3}Y^{8}Z^{11}u^{6}+\beta^{3}YZ^{18}u^{13}+\beta^{2}Y^{16}Z^{4}u^{20}$
$+\beta^{2}Y^{16}Z^{4}u^{17}+\beta^{2}Y^{9}Z^{11}u^{18}+\beta^{2}Y^{2}Z^{18}u^{13}+\beta Y^{17}Z^{4}u^{23}+\beta Y^{17}Z^{4}u^{17}$
$+\beta Y^{10}Z^{11}u^{18}+\beta Y^{3}Z^{18}u^{13}+Y^{18}Z^{4}u^{23}+Y^{18}Z^{4}u^{20}+Y^{4}Z^{18}u^{13}+Y^{4}Z^{18}u^{10},$
$a_{3}(\beta,Y,Z)$ $:=\beta^{15}Z^{6}u^{21}+\beta^{15}Z^{6}u^{15}+\beta^{15}Z^{6}u^{9}+\beta^{15}Z^{6}u^{3}+\beta^{14}YZ^{6}u^{21}+\beta^{14}YZ^{6}u^{15}$
$+\beta^{14}YZ^{6}u^{9}+\beta^{14}YZ^{6}u^{3}+\beta^{13}Y^{2}Z^{6}u^{21}+\beta^{13}Y^{2}Z^{6}u^{15}+\beta^{13}Y^{2}Z^{6}u^{9}$
$+\beta^{13}Y^{2}Z^{6}u^{3}+\beta^{11}Y^{4}Z^{6}u^{21}+\beta^{11}Y^{4}Z^{6}u^{15}+\beta^{11}Y^{4}Z^{6}u^{9}+\beta^{11}Y^{4}Z^{6}u^{3}$
$+\beta^{10}Y^{5}Z^{6}u^{21}+\beta^{10}Y^{5}Z^{6}u^{15}+\beta^{10}Y^{5}Z^{6}u^{9}+\beta^{10}Y^{5}Z^{6}u^{3}+\beta^{9}Y^{6}Z^{6}u^{15}$
$+\beta^{9}Y^{6}Z^{6}u^{3}+\beta^{8}Y^{7}Z^{6}u^{21}+\beta^{8}Y^{7}Z^{6}u^{9}+\beta^{7}Y^{8}Z^{6}u^{21}+\beta^{7}Y^{8}Z^{6}u^{15}+\beta^{7}Y^{8}Z^{6}u^{9}$
$+\beta^{7}Y^{8}Z^{6}u^{3}+\beta^{6}Y^{9}Z^{6}u^{15}+\beta^{6}Y^{9}Z^{6}u^{3}+\beta^{5}Y^{10}Z^{6}u^{21}+\beta^{5}Y^{10}Z^{6}u^{15}$
$+\beta^{5}Y^{10}Z^{6}u^{9}+\beta^{5}Y^{10}Z^{6}u^{3}+\beta^{4}Y^{11}Z^{6}u^{21}+\beta^{4}Y^{11}Z^{6}u^{9}+\beta^{3}Y^{12}Z^{6}u^{21}$
$+\beta^{3}Y^{12}Z^{6}u^{15}+\beta^{3}Y^{12}Z^{6}u^{9}+\beta^{2}Y^{13}Z^{6}u^{15}+\beta Y^{14}Z^{6}u^{15},$
$a_{4}(\beta,Y,Z)$ $:=\beta^{15}Y^{4}Zu^{18}+\beta^{15}Y^{4}Zu^{6}+\beta^{14}Y^{5}Zu^{18}+\beta^{14}Y^{5}Zu^{6}+\beta^{13}Y^{6}Zu^{18}$
$+\beta^{13}Y^{6}Zu^{6}+\beta^{12}Z^{8}u^{19}+\beta^{12}Z^{8}u^{16}+\beta^{12}Z^{8}u^{13}+\beta^{12}Z^{8}u^{10}$
$+\beta^{11}Y^{8}Zu^{18}+\beta^{11}Y^{8}Zu^{6}+\beta^{10}Y^{9}Zu^{18}+\beta^{10}Y^{9}Zu^{6}+\beta^{9}Y^{10}Zu^{18}$
$+\beta^{9}Y^{10}Zu^{6}+\beta^{9}Y^{3}Z^{8}u^{19}+\beta^{9}Y^{3}Z^{8}u^{13}+\beta^{8}Y^{4}Z^{8}u^{16}+\beta^{8}Y^{4}Z^{8}u^{10}$
$+\beta^{7}Y^{12}Zu^{18}+\beta^{7}Y^{12}Zu^{6}+\beta^{6}Y^{13}Zu^{18}+\beta^{6}Y^{13}Zu^{6}+\beta^{6}Y^{6}Z^{8}u^{19}$
$+\beta^{6}Y^{6}Z^{8}u^{10}+\beta^{5}Y^{14}Zu^{18}+\beta^{5}Y^{14}Zu^{6}+\beta^{4}Y^{8}Z^{8}u^{16}+\beta^{4}Y^{8}Z^{8}u^{13}$
$+\beta^{3}Y^{16}Zu^{18}+\beta^{3}Y^{2}Z^{15}u^{14}+\beta^{3}Y^{2}Z^{15}u^{8}+\beta^{2}Y^{17}Zu^{18}+\beta^{2}Y^{10}Z^{8}u^{19}$
$+\beta^{2}Y^{10}Z^{8}u^{10}+\beta^{2}Y^{3}Z^{15}u^{14}+\beta^{2}Y^{3}Z^{15}u^{8}+\beta Y^{18}Zu^{18}+\beta Y^{11}Z^{8}u^{19}$
$+\beta Y^{11}Z^{8}u^{13}+\beta Y^{4}Z^{15}u^{14}+\beta Y^{4}Z^{15}u^{8}+Y^{12}Z^{8}u^{19}+Y^{12}Z^{8}u^{16},$
$a_{5}(\beta,Y,Z)$ $:=\beta^{9}Z^{10}u^{11}+\beta^{9}Z^{10}u^{5}+\beta^{8}YZ^{10}u^{11}+\beta^{8}YZ^{10}u^{5}+\beta^{6}Y^{3}Z^{10}u^{11}$
$+\beta^{6}Y^{3}Z^{10}u^{5}+\beta^{4}Y^{5}Z^{10}u^{11}+\beta^{4}Y^{5}Z^{10}u^{5}+\beta^{3}Y^{6}Z^{10}u^{5}+\beta^{2}Y^{7}Z^{10}u^{11}$
$+\beta Y^{8}Z^{10}u^{11}$
$a_{6}(\beta,Y,Z)$ $:=\beta^{9}Y^{4}Z^{5}u^{20}+\beta^{9}Y^{4}Z^{5}u^{8}+\beta^{8}Y^{5}Z^{5}u^{20}+\beta^{8}Y^{5}Z^{5}u^{8}+\beta^{6}Y^{7}Z^{5}u^{20}$
$+\beta^{6}Y^{7}Z^{5}u^{8}+\beta^{6}Z^{12}u^{15}+\beta^{6}Z^{12}u^{12}+\beta^{6}Z^{12}u^{9}+\beta^{6}Z^{12}u^{6}+\beta^{4}Y^{9}Z^{5}u^{20}$
$+\beta^{4}Y^{9}Z^{5}u^{8}+\beta^{4}Y^{2}Z^{12}u^{15}+\beta^{4}Y^{2}Z^{12}u^{12}+\beta^{4}Y^{2}Z^{12}u^{9}+\beta^{4}Y^{2}Z^{12}u^{6}$
$+\beta^{3}Y^{10}Z^{5}u^{8}+\beta^{3}Y^{3}Z^{12}u^{9}+\beta^{2}Y^{11}Z^{5}u^{20}+\beta^{2}Y^{4}Z^{12}u^{15}+\beta^{2}Y^{4}Z^{12}u^{12}$
$+\beta^{2}Y^{4}Z^{12}u^{6}+\beta Y^{12}Z^{5}u^{20}+\beta Y^{5}Z^{12}u^{9}+Y^{6}Z^{12}u^{15}+Y^{6}Z^{12}u^{12},$
$a_{7}(\beta,Y,Z)$ $:=\beta^{15}Y^{2}u^{17}+\beta^{15}Y^{2}u^{5}+\beta^{14}Y^{3}u^{17}+\beta^{14}Y^{3}u^{5}+\beta^{13}Y^{4}u^{17}+\beta^{13}Y^{4}u^{5}$
$+\beta^{11}Y^{6}u^{17}+\beta^{11}Y^{6}u^{5}+\beta^{10}Y^{7}u^{17}+\beta^{10}Y^{7}u^{5}+\beta^{9}Y^{8}u^{17}+\beta^{9}Y^{8}u^{5}$
$+\beta^{7}Y^{10}u^{17}+\beta^{7}Y^{10}u^{5}+\beta^{6}Y^{11}u^{17}+\beta^{6}Y^{11}u^{5}+\beta^{5}Y^{12}u^{17}+\beta^{5}Y^{12}u^{5}$
$+\beta^{3}Y^{14}u^{17}+\beta^{3}Z^{14}u^{13}+\beta^{2}Y^{15}u^{17}+\beta^{2}YZ^{14}u^{13}+\beta Y^{16}u^{17}$
$+\beta Y^{2}Z^{14}u^{13},$
$a_{8}(\beta,Y,Z)$ $:=\beta^{6}Y^{8}Z^{2}u^{15}+\beta^{6}Y^{8}Z^{2}u^{12}+\beta^{4}Y^{10}Z^{2}u^{15}+\beta^{4}Y^{10}Z^{2}u^{12}+\beta^{3}Y^{11}Z^{2}u^{15}$
$+\beta^{3}Y^{4}Z^{9}u^{16}+\beta^{3}Y^{4}Z^{9}u^{10}+\beta^{2}Y^{12}Z^{2}u^{12}+\beta^{2}Y^{5}Z^{9}u^{16}+\beta^{2}Y^{5}Z^{9}u^{10}$
$+\beta Y^{13}Z^{2}u^{15}+\beta Y^{6}Z^{9}u^{16}+\beta Y^{6}Z^{9}u^{10},$
$a_{9}(\beta,Y,Z)$ $:=\beta^{9}Y^{2}Z^{4}u^{19}+\beta^{9}Y^{2}Z^{4}u^{7}+\beta^{8}Y^{3}Z^{4}u^{19}+\beta^{8}Y^{3}Z^{4}u^{7}+\beta^{6}Y^{5}Z^{4}u^{19}$
$+\beta^{6}Y^{5}Z^{4}u^{7}+\beta^{4}Y^{7}Z^{4}u^{19}+\beta^{4}Y^{7}Z^{4}u^{7}+\beta^{3}Y^{8}Z^{4}u^{13}+\beta^{3}Y^{8}Z^{4}u^{7}$
$+\beta^{2}Y^{9}Z^{4}u^{19}+\beta^{2}Y^{9}Z^{4}u^{13}+\beta Y^{10}Z^{4}u^{19}+\beta Y^{10}Z^{4}u^{13},$
$a_{{10}}(\beta,Y,Z)$ $:=\beta^{3}Y^{5}Z^{6}u^{17}+\beta^{2}Y^{6}Z^{6}u^{17}+\beta Y^{7}Z^{6}u^{17}+Y^{8}Z^{6}u^{17}+Y^{8}Z^{6}u^{14},$
$a_{{11}}(\beta,Y,Z)$ $:=\beta^{3}Y^{2}Z^{8}u^{9}+\beta^{2}Y^{3}Z^{8}u^{9}+\beta Y^{4}Z^{8}u^{9},$
$a_{{12}}(\beta,Y,Z)$ $:=\beta^{9}Z^{3}u^{18}+\beta^{9}Z^{3}u^{12}+\beta^{8}YZ^{3}u^{18}+\beta^{8}YZ^{3}u^{12}+\beta^{6}Y^{3}Z^{3}u^{18}+\beta^{6}Y^{3}Z^{3}u^{12}$
$+\beta^{4}Y^{5}Z^{3}u^{18}+\beta^{4}Y^{5}Z^{3}u^{12}+\beta^{3}Y^{6}Z^{3}u^{18}+\beta^{2}Y^{7}Z^{3}u^{12}+\beta Y^{8}Z^{3}u^{12},$
$a_{{13}}(\beta,Y,Z)$ $:=0,$
$a_{{14}}(\beta,Y,Z)$ $:=\beta^{9}Yu^{19}+\beta^{9}Yu^{13}+\beta^{8}Y^{2}u^{19}+\beta^{8}Y^{2}u^{13}+\beta^{6}Y^{4}u^{19}+\beta^{6}Y^{4}u^{10}$
$+\beta^{4}Y^{6}u^{19}+\beta^{4}Y^{6}u^{10}+\beta^{3}Y^{7}u^{13}+\beta^{3}Z^{7}u^{14}+\beta^{2}Y^{8}u^{19}+\beta^{2}Y^{8}u^{13}$
$+\beta^{2}Y^{8}u^{10}+\beta^{2}YZ^{7}u^{14}+\beta Y^{9}u^{19}+\beta Y^{2}Z^{7}u^{14},$
$a_{{15}}(\beta,Y,Z)$ $:=\beta^{3}Y^{4}Z^{2}u^{17}+\beta^{3}Y^{4}Z^{2}u^{11}+\beta^{2}Y^{5}Z^{2}u^{17}+\beta^{2}Y^{5}Z^{2}u^{11}+\beta Y^{6}Z^{2}u^{17}+\beta Y^{6}Z^{2}u^{11},$
$a_{16}(\beta,Y,Z)$ $:=\beta^{3}YZ^{4}u^{15}+\beta^{2}Y^{2}Z^{4}u^{15}+\beta Y^{3}Z^{4}u^{15}+Y^{4}Z^{4}u^{15}+Y^{4}Z^{4}u^{12},$
$a_{{17}}(\beta,Y,Z)$ $:=0,$
$a_{{18}}(\beta,Y,Z)$ $:=\beta^{3}Y^{2}Zu^{16}+\beta^{2}Y^{3}Zu^{16}+\beta Y^{4}Zu^{16},$
$a_{{19}}(\beta,Y,Z)$ $:=0,$
$a_{{20}}(\beta,Y,Z)$ $:=0,$
$a_{{21}}(\beta,Y,Z)$ $:=\beta^{3}u^{15}+\beta^{2}Yu^{15}+\beta Y^{2}u^{15}.$
Proposition 4.
Each element of
$\Theta:=$$\{(x,y,z,a,b,c)\in\mathbb{F}_{q}^{6}:(b+y)h(b,x,y,z)\neq 0,\ c=f(a,b,x,y,z),$
$\qquad\qquad\qquad\qquad a=g(b,x,y,z)/h(b,x,y,z),\sum_{i=0}^{21}a_{i}(b,y,z)x^{i}=0\}$
is a solution of System (2).
Proof.
This is easily checked via direct computations in MAGMA [6]. In particular, from the first equation in System (2) one gets $\gamma=f(\alpha,\beta,X,Y,Z)$, and combining this condition with the second equation one can check that $\alpha=g(\beta,X,Y,Z)/h(\beta,X,Y,Z)$ holds. Finally, let $H(\beta,X,Y,Z)$ be the polynomial obtained by combining the third equation in System (2) with $\gamma=f(\alpha,\beta,X,Y,Z)$ and $\alpha=g(\beta,X,Y,Z)/h(\beta,X,Y,Z)$, after clearing the denominators. It can be checked that the polynomial $\sum_{i=0}^{21}a_{i}(\beta,Y,Z)X^{i}$ is a factor of $H(\beta,X,Y,Z)$, thus completing the proof. ∎
Proposition 5.
Let $\mathcal{V}\subset\mathbb{A}^{6}(\mathbb{F}_{q})$ be the variety defined by
$\begin{cases}X_{3}=f(X_{1},X_{2},X_{4},X_{5},X_{6})\\ X_{1}=g(X_{2},X_{4},X_{5},X_{6})/h(X_{2},X_{4},X_{5},X_{6})\\ \sum_{i=0}^{21}a_{i}(X_{2},X_{5},X_{6})X_{4}^{i}=0.\end{cases}$
Then $\mathcal{V}$ contains an absolutely irreducible component defined over $\mathbb{F}_{q}$, distinct from $X_{1}=X_{2}=X_{3}=0$, of degree at most $27$, and not contained in $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$.
Proof.
Since $\sum_{i=0}^{21}a_{i}(X_{2},X_{5},X_{6})X_{4}^{i}$ is a homogeneous polynomial of degree $24$ it defines a surface $\mathcal{S}$ in $\mathbb{P}^{3}(\mathbb{F}_{q})$. First note that $X_{2}$ is not a factor of $\sum_{i=0}^{21}a_{i}(X_{2},X_{5},X_{6})X_{4}^{i}$, otherwise $u\in\mathbb{F}_{4}^{*}$, a contradiction. Also, $\mathcal{S}$ contains an absolutely irreducible non-repeated component defined over $\mathbb{F}_{q}$, since its intersection with the plane $X_{4}=0$ is $a_{0}(X_{2},X_{5},X_{6})=0$ and
$X_{2}\mid a_{0}(X_{2},X_{5},X_{6})\textrm{ and }X_{2}^{2}\nmid a_{0}(X_{2},X_{5},X_{6}),$
by Lemma 2 there exists an absolutely irreducible component $\mathcal{S}^{\prime}$ defined over $\mathbb{F}_{q}$ whose intersection with $X_{4}=0$ is $X_{2}=0$.
Since the intersection between $X_{4}=0$ and $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$ is
$u^{5}X_{6}^{2}(X_{2}+X_{5})^{3}((u^{2}+u^{5})X_{2}^{3}X_{5}^{4}+(u^{2}+u^{5})X_{2}^{2}X_{5}^{5}+(u^{2}+u^{5})X_{2}X_{5}^{6}+u^{5}X_{5}^{7}+X_{6}^{7})=0,$
$\mathcal{S}^{\prime}$ cannot be contained in $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$.
Clearly, $\mathcal{S}^{\prime}$ extends to a variety $\mathcal{V}^{\prime}$ contained in $\mathcal{V}$, absolutely irreducible and $\mathbb{F}_{q}$-rational (by considering the two extra equations $X_{3}=f(X_{1},X_{2},X_{4},X_{5},X_{6})$ and $X_{1}=g(X_{2},X_{4},X_{5},X_{6})/h(X_{2},X_{4},X_{5},X_{6}))$. Since $\mathcal{S}^{\prime}$ is different from $X_{2}=0$, $\mathcal{V}^{\prime}$ is different from $X_{1}=X_{2}=X_{3}=0$. The degree of $\mathcal{V}^{\prime}$ is upper bounded by the degree of $\mathcal{V}$, that is at most $273^{3}=$ 27. ∎
The following theorem is the main result of our paper.
Theorem 6.
If $m\geq 25$, then Conjecture 3 is true.
Proof.
By Proposition 5 and Theorem 1, if $m$ is large enough, the absolutely irreducible $\mathbb{F}_{q}$-rational component $\mathcal{V}^{\prime}$ (and thus $\mathcal{V}$) possesses roughly $q^{3}$ of $\mathbb{F}_{q}$-rational points. Since $\mathcal{V}^{\prime}$ is not $X_{1}=X_{2}=X_{3}=0$ and it not contained in $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$, the set $\Theta$ is not empty and the validity of the conjecture follows.
The bound $m\geq 25$ can be shown via Theorem 1. First, notice that points of $\mathcal{V}^{\prime}$ not belonging to $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$ correspond to quadruples $(x_{2},x_{4},x_{5},x_{6})\in\mathbb{F}_{q}^{4}$ such that $\sum_{i=0}^{21}a_{i}(x_{2},x_{4},x_{5})x_{4}^{i}=0$ and $(x_{2}+x_{5})h(x_{2},x_{4},x_{5},x_{6})\neq 0$. It can be easily verified that the intersections of $\mathcal{S}$ and $(X_{2}+X_{5})h(X_{2},X_{4},X_{5},X_{6})=0$ with $X_{4}=0$ are coprime: this shows that they do not share any component. Also, since the degree of these surfaces is $24$ and $12$ respectively, their intersection is a projective space curve of degree $288$ and it contains at most $288q$ projective $\mathbb{F}_{q}$-rational points; see [11]. This shows the existence of at most $288q(q-1)$ nonzero quadruples $(x_{1},x_{2},x_{3},x_{4})\in\mathbb{F}_{q}^{4}$ such that $\sum_{i=0}^{21}a_{i}(x_{2},x_{4},x_{5})x_{4}^{i}=0$ and $(x_{2}+x_{5})h(x_{2},x_{4},x_{5},x_{6})=0$. Arguing similarly, there are at most $24q(q-1)$ nonzero quadruples $(x_{1},x_{2},x_{3},x_{4})\in\mathbb{F}_{q}^{4}$ satisfying $\sum_{i=0}^{21}a_{i}(x_{2},x_{4},x_{5})x_{4}^{i}=0$ and $x_{2}=0$ (and thus lying on $X_{1}=X_{2}=X_{3}=0$). By Theorem 1 the number of $\mathbb{F}_{q}$-rational points in $\Theta$ is at least
$q^{3}-25\cdot 26\cdot q^{5/2}-5\cdot(27)^{13/3}q^{2}-312q(q-1),$
that is larger than 00 when $m\geq 25$ and thus Conjecture 3 holds true. ∎
3 Conclusion
In this paper, using algebraic geometry tools, we show a conjecture of Beierle, Carlet, Leander, Perrin from 2022, when the underlying field has dimension $m\geq 25$.
Acknowledgements
The third-named author (PS) would like to thank the first-named author (DB) for the invitation at the Dipartimento di Matematica e Informatica at Università degli Studi di Perugia, Italy, and the great working conditions while this paper was being written. The first-named author (DB) thanks the Italian National Group for Algebraic and Geometric Structures and their Applications (GNSAGA—INdAM) which supported the research.